If you throw candies to the people behind the float in the opposite direction at a speed of 1 m/s, how fast will the candy reach the people?


1
Expert's answer
2021-06-23T13:09:21-0400

Gives

Speed(u)=1m/sec

We know that

Suppose candies through and t time after people obtained candies

"v=u+at"

a=g

"v=u+gt"

"v=1+9.8t"

Put t value and find v value

the candy reach the people of speed

"v=1+9.8t"